$SPAD/src/input heugcd.input

Waldek Hebisch and Timothy Daly

December 3, 2016

Abstract

This is a test of revision 256 of FriCAS which changed the gcd function in listgcd.spad.pamphlet.

Contents

2

Of these examples, all but the first used to return 1. This is a test of revision 256 of FriCAS which changed the gcd function in listgcd.spad.pamphlet.

— * —

)set break resume
)spool heugcd.output
)set message test on
)set message auto off
)clear all
--S 1 of 5
gcd([0,0,x^2-1,x^2+2*x+1])
--R
--R   (1)  x + 1
--R                                                    Type: Polynomial(Integer)
--E 1
--S 2 of 5
gcd([0,0,x^2-1,x^2+2*x+1])$HeuGcd(SparseUnivariatePolynomial Integer)
--R
--R   (2)  ? + 1
--R                                    Type: SparseUnivariatePolynomial(Integer)
--E 2
--S 3 of 5
gcd(6*x^2-1,36*x^2+12*x+1)
--R
--R   (3)  1
--R                                                    Type: Polynomial(Integer)
--E 3
--S 4 of 5
gcd([36*x^2-1,36*x^2+12*x+1])
--R
--R   (4)  6x + 1
--R                                                    Type: Polynomial(Integer)
--E 4
--S 5 of 5
gcd([36*x^2-1,36*x^2+12*x+1])$HeuGcd(SparseUnivariatePolynomial Integer)
--R
--R   (5)  6? + 1
--R                                    Type: SparseUnivariatePolynomial(Integer)
--E 5
)spool
)lisp (bye)

———-

References

[1] http://fricas.svn.sourceforge.net/fricas/?rev=256&view=rev